Use mathematical induction to prove that for all n ∈ Z+
5 + 22 + 39 + · · · + (17n - 12) = n ·(17n - 7)/2
4)(20) The relation R: Z x Z is defined as
for a, b ∈ Z, (a, b) ∈ R if a + b is even.
Prove all the properties: reflexive, symmetric, anti-symmetric, transitive
that relation R has. If R does not have any of these properties, explain why.
Is R an equivalence relation?
Answer #1

112 5+ es +39€ -+ (170-12) = 2 (199-7) 2 - tno zt we prove by Induction DEL P(1):5= +(1:6-7) -5 2 Lence ph is true for n=1 re(a, b ER if a+b es eren fa, bez (1) Reflesive :- aGz hence (a, a) ER tuen ata=2a even R is reflenire. ☺ Symmetic:- (a, b) ER